# TUTORIAL ARTICLE Contraction Theory for Nonlinear Stability Analysis and Learning-based Control: A Tutorial Overview

### v7

@@ 3.1.2 Linearization of Nonlinear Systems
If a contraction metric of a given nonlinear system is unknown, we could linearize it to apply methodologies inspired by LTV systems theory such as $\mathcal{H}_{\infty}$ control Doyle et al. [1989], Khargonekar and Rotea [1991], Xie and de Souza Carlos [1992], Scherer et al. [1997], Hinrichsen and Pritchard [1998], Chen and Zhang [2004], iterative Linear Quadratic Regulator (iLQR) Li and Todorov [2004, 2006], or Extended Kalman Filter (EKF). Their stability is typically analyzed by decomposing $f(x,t)$ as f(x,t)=Ax+(f(x,t)Ax)f(x,t)=Ax+(f(x,t)-Ax) assuming that the nonlinear part f(x,t)Axf(x,t)-Ax is bounded, or by finding a local contraction region for the sake of local exponential stability as in Bonnabel and Slotine [2015], Tsukamoto and Chung [2021]. Since the decomposition f(x,t)=Ax+(f(x,t)Ax)f(x,t)=Ax+(f(x,t)-Ax) allows applying the result of Theorem 2.4, we could exploit the techniques in Sec. 3 and Sec. 4 for providing formal robustness and optimality guarantees for the LTV systems-type approaches. For systems whose nonlinear part f(x,t)Axf(x,t)-Ax is not necessarily bounded, Sec. 9.1 elucidates how contraction theory can be used to stabilize them with the learned residual dynamics for control synthesis.
3.1.3 State-Dependent Coefficient (SDC) Formulation
@@
It is shown in Tsukamoto and Chung [2021], Dani et al. [2015], Tsukamoto et al. [2021], Tsukamoto and Chung [2021, 2019] that the SDC-based control and estimation Cloutier [1997], Banks et al. [2007], Chang and Chung [2009], Çimen [2012], which capture nonlinearity using a state-dependent matrix $A(x,t)$ s.t. $f(x,t)=A(x,t)x$ (e.g., we have $A(x,t)=\cos x$ for $f(x,t)=x\cos x$), result in exponential boundedness of system trajectories both for deterministic and stochastic systems due to Theorems 2.4 and 2.5 Tsukamoto and Chung [2021]. Because of the extended linear form of SDC (see Table 3), the results to be presented in Sec. 34 based on the SDC formulation are applicable to linearized dynamics that can be viewed as an LTV system with some modifications (see Remark 3.2).
@@
This idea is slightly generalized in Tsukamoto et al. [2021] to explicitly consider incremental stability with respect to a target trajectory (e.g., $x_{d}$ for control and $x$ for estimation) instead of using $A(x,t)x=f(x,t)$. Let us derive the following lemma for this purpose Tsukamoto et al. [2021], Çimen [2012], Tsukamoto and Chung [2021], Dani et al. [2015], Tsukamoto and Chung [2019].
@@
Let us illustrate how Lemma 3.1 can be used in practice taking the following nonlinear system as an example:
@@
The major advantage of the formalism in Lemma 3.1 lies in its systematic connection to LTV systems based on uniform controllability and observability, adequately accounting for the nonlinear nature of underlying dynamics through $A(\varrho,x,x_{d},u_{d},t)$ for global stability, as shall be seen in Sec. 3 and Sec. 4. Since $A$ depends also on $(x_{d},u_{d})$ in this case unlike the original SDC matrix, we could consider contraction metrics using a positive definite matrix $M(x,x_{d},u_{d},t)$ instead of $M(x,t)$ in Definition 2.3, to improve the representation power of $M$ at the expense of computational efficiency. Another interesting point is that the non-uniqueness of $A$ in Lemma 3.1 for $n\geq 2$ creates additional degrees of freedom for selecting the coefficients $\varrho$, which can also be treated as decision variables in constructing optimal contraction metrics as proposed in Tsukamoto and Chung [2021], Dani et al. [2015], Tsukamoto and Chung [2019].
We focus mostly on the generalized SDC formulation in Sec. 3 and Sec. 4, as it yields optimal control and estimation laws with global stability Tsukamoto et al. [2021] while keeping the analysis simple enough to be understood as in LTV systems theory.
@@
This does not mean that contraction theory works only for the SDC parameterized nonlinear systems but implies that it can be used with the other techniques discussed in Sec. 3.1. For example, due to the extended linear form given in Table 3, the results to be presented in Sec. 3 and in Sec. 4 based on the SDC formulation are applicable to linearized dynamics that can be viewed as an LTV system with some modifications, regarding the dynamics modeling error term as an external disturbance as in Sec. 3.1.2. Also, the original SDC formulation with respect to a fixed point (e.g., $(s,\bar{s},\bar{u})=(x,0,0)$ in Lemma 3.1) can still be used to obtain contraction conditions independent of a target trajectory $(x_{d},u_{d})$ (see Theorem 3.2 for details).
3.1.4 Control Contraction Metric (CCM) Formulation
We could also consider using the partial derivative of $f$ of the dynamical system directly for control synthesis through differential state feedback $\delta u=k(x,\delta x,u,t)$. This idea, formulated as the concept of a CCM Manchester and Slotine [2017], Singh et al. [2017, 2020], Wang et al. [2019, 2020], Manchester and Slotine [2018], constructs contraction metrics with global stability guarantees independently of target trajectories, achieving greater generality while requiring added computation in evaluating integrals involving minimizing geodesics. Similar to the CCM, we could design a state estimator using a general formulation based on geodesics distances between trajectories Pham and Slotine [2013], Manchester [2018]. These approaches are well compatible with the convex optimization-based schemes in Sec. 4, and hence will be discussed in Sec. 4.3.
@@ 3.1.4 Control Contraction Metric (CCM) Formulation
For control non-affine nonlinear systems, we could find $f(x,u,t)-f(x_{d},u_{d},t)=A(x,x_{d},u,u_{d},t)(x-x_{d})+B(x,x_{d},u,u_{d},t)(u-u_{d})\allowbreak$ by Lemma 3.1 on the SDC formulation and use it in Theorem 4.2, although (3.15) has to be solved implicitly as $B$ depends on $u$ in this case. A similar approach for the CCM formulation can be found in Wang et al. [2020, 2019]. As discussed in Example 3.2, designing such implicit control laws will be discussed in Lemma 9.1 and Theorem 9.2 of Sec. 9.1.
3.2 LMI Conditions for Contraction Metrics
We design a nonlinear feedback tracking control law parameterized by a matrix-valued function $M(x,x_{d},u_{d},t)$ (or $M(x,t)$, see Theorem 3.2) as follows:
@@ 3.2 LMI Conditions for Contraction Metrics
where $R(x,x_{d},u_{d},t)\succ 0$ is a weight matrix on the input $u$ and $M(x,x_{d},u_{d},t)\succ 0$ is a positive definite matrix (which satisfies the matrix inequality constraints for a contraction metric, to be given in Theorem 3.1). As discussed in Sec. 3.1.3, the extended linear form of the tracking control (3.15) enables LTV systems-type approaches to Lyapunov function construction, while being general enough to capture the nonlinearity of the underlying dynamics due to Lemma 3.2 Tsukamoto and Chung [2021].
@@
gives the desired relation Tsukamoto and Chung [2021]. $\qed$
@@
Lemma 3.2 implies that designing optimal $k$ of $u=k(x,x_{d},u_{d},t)$ reduces to designing the optimal gain $K(x,x_{d},u_{d},t)$ of $u=u_{d}-K(x,x_{d},u_{d},t)(x-x_{d})$. We could also generalize this idea further using the CCM-based differential feedback controller $\delta u=k(x,\delta x,u,t)$ Manchester and Slotine [2017], Singh et al. [2017, 2020], Wang et al. [2019, 2020], Manchester and Slotine [2018] (see Theorem 4.6).
Substituting (3.15) into (3.1) and (3.2) yields the following virtual system of a smooth path $q(\mu,t)$, parameterized by $\mu\in[0,1]$ to have $q(0,t)=x_{d}$ and $q(1,t)=x$, for partial contraction in Theorem 2.2:
@@
In Tsukamoto and Chung [2021], Tsukamoto et al. [2021], Tsukamoto and Chung [2021, 2019], it is proposed that the contraction conditions of Theorems 2.1 and 2.5 for the closed-loop dynamics (3.17) and (3.18) can be expressed as convex constraints as summarized in Theorem 3.1.
@@
Let $\beta$ be defined as $\beta=0$ for deterministic systems (3.1) and
@@
for stochastic systems (3.2), respectively, where $\bar{g}_{c}$ is given in (3.2), $L_{m}$ is the Lipschitz constant of $\partial M/\partial x_{i}$ for $M$ of (3.15), and $\alpha_{G}\in\mathbb{R}_{>0}$ is an arbitrary constant as in Theorem 2.5. Also, let $W=M(x,x_{d},u_{d},t)^{-1}$ (or $W=M(x,t)^{-1}$, see Theorem 3.2), $\bar{W}=\nu W$, and $\nu=\overline{m}$. Then the following three matrix inequalities are equivalent:
@@
Note that $\nu$ and $\bar{W}$ are required for (3.24) and (3.25) and the arguments $(x,x_{d},u_{d},t)$ for each matrix are suppressed for notational simplicity.
Furthermore, under these equivalent contraction conditions, Theorems 2.4 and 2.5 hold for the virtual systems (3.17) and (3.18), respectively. In particular, if $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ of (2.32) holds, or equivalently
@@
where $V_{s\ell}=\int^{x}_{x_{d}}\delta q^{\top}M\delta q$ and $V_{\ell}=\int^{x}_{x_{d}}\|\Theta\delta q\|$ are as given in Theorem 2.3 with $M=\Theta^{\top}\Theta$, the disturbance bounds $\bar{d}_{c}$ and $\bar{g}_{c}$ are given in (3.1) and (3.2), respectively, and $C_{C}=\bar{g}_{c}^{2}({2}{\alpha_{G}}^{-1}+1)$. Note that for stochastic systems, the probability that $\|x-x_{d}\|$ is greater than or equal to $\varepsilon\in\mathbb{R}_{>0}$ is given as
@@
Substituting (3.19) into (3.22) gives (3.23). Since $\nu>0$ and $W\succ 0$, multiplying (3.23) by $\nu$ and then by $W$ from both sides preserves matrix definiteness. Also, the resultant inequalities are equivalent to the original ones [Boyd et al., 1994, p. 114]. These operations performed on (3.23) yield (3.24). If $\beta=\alpha_{s}>0$ for stochastic systems, applying Schur’s complement lemma [Boyd et al., 1994, p. 7] to (3.24) results in the Linear Matrix Inequality (LMI) constraint (3.25) in terms of $\bar{W}$ and $\nu$. Therefore, (3.22)–(3.25) are indeed equivalent.
Also, since we have $\|\partial_{\mu}d(\mu,x,t)\|\leq\bar{d}_{c}$ for $d$ in (3.17) and $\|\partial_{\mu}G(\mu,x,t)\|_{F}^{2}\leq\bar{g}_{c}^{2}$ for $G$ in (3.18), the virtual systems in (3.17) and (3.18) clearly satisfy the conditions of Theorems 2.4 and 2.5 if it is equipped with (3.22), which is equivalent to (3.23)–(3.25). This implies the exponential bounds (3.27)–(3.29) rewritten using $\chi=\overline{m}/\underline{m}$, following the proofs of Theorems 2.4 and 2.5. $\qed$
@@
Although the conditions (3.22)–(3.25) depend on $(x_{d},u_{d})$, we could also use the SDC formulation with respect to a fixed point Tsukamoto and Chung [2021, 2019] in Lemma 3.1 to make them independent of the target trajectory as in the following theorem.
@@
Let $(\bar{x},\bar{u})$ be a fixed point selected arbitrarily in $\mathbb{R}^{n}\times\mathbb{R}^{m}$, e.g., $(\bar{x},\bar{u})=(0,0)$, and let $A(x,t)$ be an SDC matrix constructed with $({s},\bar{s},\bar{u})=(x,\bar{x},\bar{u})$ in Lemma 3.1, i.e.,
@@
Suppose that the contraction metric of Theorem 3.1 is designed by $M(x,t)$ with $A$ of (3.30), independently of the target trajectory $(x_{d},u_{d})$, and that the systems (3.1) and (3.2) are controlled by
@@
with such $M(x,t)$, where $R(x,t)\succ 0$ is a weight matrix on $u$. If the function $\phi(x,x_{d},u_{d},t)=A(\varrho,x,t)(x_{d}-\bar{x})+B(x,t)(u_{d}-\bar{u})$ is Lipschitz in $x$ with its Lipschitz constant $\bar{L}$, then Theorem 3.1 still holds with $\alpha$ of the conditions (3.22)–(3.25) replaced by $\alpha+\bar{L}\sqrt{\overline{m}/\underline{m}}$. The same argument holds for state estimation of Theorem 4.3 to be discussed in Sec. 4.2.
@@
where $K(x,t)=R(x,t)^{-1}B(x,t)^{\top}M(x,t)$. Following the proof of Theorem 3.1, the computation of $\dot{V}$, where $V=\delta q^{\top}M(x,t)\delta q$, yields an extra term
@@
As demonstrated in Tsukamoto and Chung [2021], we could directly use the extra term $2\delta q^{\top}M({\partial\phi}/{\partial q})\delta q$ of (3.33) in (3.22)–(3.25) without upper-bounding it, although now the constraints of Theorem 3.2 depend on $(x,q,t)$ instead of $(x,t)$. Also, the following two inequalities given in Tsukamoto and Chung [2021] with $\bar{\gamma}=\nu\gamma$, $\gamma\in\mathbb{R}_{\geq 0}$:
@@
are combined as one LMI (3.25) in Theorems 3.1 and 3.2.
@@
The inequalities in Theorem 3.1 can be interpreted as in the Riccati inequality in $\mathcal{H}_{\infty}$ control. Consider the following system:
@@
where $A\in\mathbb{R}^{n\times n}$, $B_{u}\in\mathbb{R}^{n\times m}$, $B_{w}\in\mathbb{R}^{n\times w}$, and $C_{z}\in\mathbb{R}^{o\times n}$ are constant matrices, $w\in\mathbb{R}^{w}$ is an exogenous input, and $z\in\mathbb{R}^{o}$ is a system output. As shown in Van der Schaft [1992] and [Boyd et al., 1994, p. 109], there exists a state feedback gain $K=R^{-1}B_{u}^{\top}P$ such that the $\mathcal{L}_{2}$ gain of the closed-loop system (3.35), $\sup_{\|w\|\neq 0}{\|z\|}/{\|w\|}$, is less than or equal to $\gamma$ if
@@
has a solution $P\succ 0$, where $R\succ 0$ is a constant weight matrix on the input $u$. If we select $B_{w}$ and $C_{z}$ to have $B_{w}B_{w}^{\top}\succeq(P^{-1})^{2}$ and $C_{z}^{\top}C_{z}\succeq 2\alpha P$ for some $\alpha>0$, the contraction condition (3.23) in Theorem 3.1 can be satisfied with $M=P$, $B=B_{u}$, and $\beta=1/\gamma^{2}$ due to (3.36).
In Sections 3.3 and 3.4, we will discuss the relationship to input-output stability theory as in Example 3.4, using the results of Theorem 3.1.
3.3 Bounded Real Lemma in Contraction Theory
The LMI (3.25) of Theorem 3.1 can be interpreted as the LMI condition for the bounded real lemma [Başar and Bernhard, 2008, p. 369]. Let us consider the following Hurwitz linear system with $\varsigma(t)\in\mathcal{L}_{2e}$, (i.e., $\|(\varsigma)_{\tau}\|_{\mathcal{L}_{2}}<\infty$ for $\tau\in\mathbb{R}_{\geq 0}$, see Sec. 1.3):

### v8

@@ 3.1.2. Linearization of Nonlinear Systems
If a contraction metric of a given nonlinear system is unknown, we could linearize it to apply methodologies inspired by LTV systems theory such as $\mathcal{H}_{\infty}$ control 29425, 85062, 151101, 599969, doi:10.1137/S0363012996301336, 1259458, iterative Linear Quadratic Regulator (iLQR) ilqr, silqr, or Extended Kalman Filter (EKF). Their stability is typically analyzed by decomposing $f(x,t)$ as f(x,t)=Ax+(f(x,t)Ax)f(x,t)=Ax+(f(x,t)-Ax) assuming that the nonlinear part f(x,t)Axf(x,t)-Ax is bounded, or by finding a local contraction region for the sake of local exponential stability as in 6849943, ncm. Since the decomposition f(x,t)=Ax+(f(x,t)Ax)f(x,t)=Ax+(f(x,t)-Ax) allows applying the result of Theorem LABEL:THM:Thm:Robust_contraction_original, we could exploit the techniques in Sec. 3 and Sec. 4 for providing formal robustness and optimality guarantees for the LTV systems-type approaches. For systems whose nonlinear part f(x,t)Axf(x,t)-Ax is not necessarily bounded, Sec. 9.1 elucidates how contraction theory can be used to stabilize them with the learned residual dynamics for control synthesis.
3.1.3. State-Dependent Coefficient (SDC) Formulation
@@
It is shown in mypaperTAC, Ref:Stochastic, nscm, ncm, mypaper that the SDC-based control and estimation sddre, Banks2007, chang2009exponential, survey_SDRE, which capture nonlinearity using a state-dependent matrix $A(x,t)$ s.t. $f(x,t)=A(x,t)x$ (e.g., we have $A(x,t)=\cos x$ for $f(x,t)=x\cos x$), result in exponential boundedness of system trajectories both for deterministic and stochastic systems due to Theorems LABEL:THM:Thm:Robust_contraction_original and LABEL:THM:Thm:robuststochastic ncm. Because of the extended linear form of SDC (see Table 3), the results to be presented in Sec. 34 based on the SDC formulation are applicable to linearized dynamics that can be viewed as an LTV system with some modifications (see Remark 3.2).
@@
This idea is slightly generalized in nscm to explicitly consider incremental stability with respect to a target trajectory (e.g., $x_{d}$ for control and $x$ for estimation) instead of using $A(x,t)x=f(x,t)$. Let us derive the following lemma for this purpose nscm, survey_SDRE, mypaperTAC, Ref:Stochastic, mypaper.
@@
SDC Formulation Let us illustrate how Lemma 3.1 can be used in practice, taking the following nonlinear system as an example:
@@
The major advantage of the formalism in Lemma 3.1 lies in its systematic connection to LTV systems based on uniform controllability and observability, adequately accounting for the nonlinear nature of underlying dynamics through $A(\varrho,x,x_{d},u_{d},t)$ for global stability, as shall be seen in Sec. 3 and Sec. 4. Since $A$ depends also on $(x_{d},u_{d})$ in this case unlike the original SDC matrix, we could consider contraction metrics using a positive definite matrix $M(x,x_{d},u_{d},t)$ instead of $M(x,t)$ in Definition LABEL:DEF:Def:contraction, to improve the representation power of $M$ at the expense of computational efficiency. Another interesting point is that the non-uniqueness of $A$ in Lemma 3.1 for $n\geq 2$ creates additional degrees of freedom for selecting the coefficients $\varrho$, which can also be treated as decision variables in constructing optimal contraction metrics as proposed in mypaperTAC, Ref:Stochastic, mypaper.
We focus mostly on the generalized SDC formulation in Sec. 3 and Sec. 4, as it yields optimal control and estimation laws with global stability nscm while keeping the analysis simple enough to be understood as in LTV systems theory.
@@
This does not mean that contraction theory works only for the SDC parameterized nonlinear systems but implies that it can be used with the other techniques discussed in Sec. 3.1. For example, due to the extended linear form given in Table 3, the results to be presented in Sec. 3 and in Sec. 4 based on the SDC formulation are applicable to linearized dynamics that can be viewed as an LTV system with some modifications, regarding the dynamics modeling error term as an external disturbance as in Sec. 3.1.2. Also, the original SDC formulation with respect to a fixed point (e.g., $(s,\bar{s},\bar{u})=(x,0,0)$ in Lemma 3.1) can still be used to obtain contraction conditions independent of a target trajectory $(x_{d},u_{d})$ (see Theorem LABEL:THM:Thm:fixed_sdc for details).
3.1.4. Control Contraction Metric (CCM) Formulation
We could also consider using the partial derivative of $f$ of the dynamical system directly for control synthesis through differential state feedback $\delta u=k(x,\delta x,u,t)$. This idea, formulated as the concept of a CCM ccm, 7989693, 47710, WANG201944, vccm, rccm, constructs contraction metrics with global stability guarantees independently of target trajectories, achieving greater generality while requiring added computation in evaluating integrals involving minimizing geodesics. Similar to the CCM, we could design a state estimator using a general formulation based on geodesics distances between trajectories scrm, estimation_ccm. These approaches are well compatible with the convex optimization-based schemes in Sec. 4, and hence will be discussed in Sec. 4.3.
@@ 3.1.4. Control Contraction Metric (CCM) Formulation
For control non-affine nonlinear systems, we could find $f(x,u,t)-f(x_{d},u_{d},t)=A(x,x_{d},u,u_{d},t)(x-x_{d})+B(x,x_{d},u,u_{d},t)(u-u_{d})\allowbreak$ by Lemma 3.1 on the SDC formulation and use it in Theorem LABEL:THM:Thm:CV-STEM, although (3.15) has to be solved implicitly as $B$ depends on $u$ in this case. A similar approach for the CCM formulation can be found in vccm, WANG201944. As discussed in Example LABEL:EX:ex:nonaffine2, designing such implicit control laws will be discussed in Lemma 9.1 and Theorem LABEL:THM:Thm:neurallander of Sec. 9.1.
3.2. LMI Conditions for Contraction Metrics
We design a nonlinear feedback tracking control law parameterized by a matrix-valued function $M(x,x_{d},u_{d},t)$ (or $M(x,t)$, see Theorem LABEL:THM:Thm:fixed_sdc) as follows:
@@ 3.2. LMI Conditions for Contraction Metrics
where $R(x,x_{d},u_{d},t)\succ 0$ is a weight matrix on the input $u$ and $M(x,x_{d},u_{d},t)\succ 0$ is a positive definite matrix (which satisfies the matrix inequality conditions for a contraction metric, to be given in Theorem LABEL:THM:Thm:CVSTEM:LMI). As discussed in Sec. 3.1.3, the extended linear form of the tracking control (3.15) enables LTV systems-type approaches to Lyapunov function construction, while being general enough to capture the nonlinearity of the underlying dynamics due to Lemma 3.2 lagros.
@@
gives the desired relation lagros. $\qed$
@@
Lemma 3.2 implies that designing optimal $k$ of $u=k(x,x_{d},u_{d},t)$ reduces to designing the optimal gain $K(x,x_{d},u_{d},t)$ of $u=u_{d}-K(x,x_{d},u_{d},t)(x-x_{d})$. We could also generalize this idea further using the CCM-based differential feedback controller $\delta u=k(x,\delta x,u,t)$ ccm, 7989693, 47710, WANG201944, vccm, rccm (see Theorem LABEL:THM:Thm:ccm_cvstem).
Substituting (3.15) into (3.1) and (3.2) yields the following virtual system of a smooth path $q(\mu,t)$, parameterized by $\mu\in[0,1]$ to have $q(0,t)=x_{d}$ and $q(1,t)=x$, for partial contraction in Theorem LABEL:THM:Thm:partial_contraction:
@@
In mypaperTAC, nscm, ncm, mypaper, it is proposed that the contraction conditions of Theorems LABEL:THM:Thm:contraction and LABEL:THM:Thm:robuststochastic for the closed-loop dynamics (3.17) and (3.18) can be expressed as convex conditions as summarized in Theorem LABEL:THM:Thm:CVSTEM:LMI.
@@
Convexity of Contraction Conditions for ControlThm:CVSTEM:LMI Let $\beta$ be defined as $\beta=0$ for deterministic systems (3.1) and
@@
for stochastic systems (3.2), respectively, where $\bar{g}_{c}$ is given in (3.2), $L_{m}$ is the Lipschitz constant of $\partial M/\partial x_{i}$ for $M$ of (3.15), and $\alpha_{G}\in\mathbb{R}_{>0}$ is an arbitrary constant as in Theorem LABEL:THM:Thm:robuststochastic. Also, let $W=M(x,x_{d},u_{d},t)^{-1}$ (or $W=M(x,t)^{-1}$, see Theorem LABEL:THM:Thm:fixed_sdc), $\bar{W}=\nu W$, and $\nu=\overline{m}$. Then the following three matrix inequalities are equivalent:
@@
Note that $\nu$ and $\bar{W}$ are required for (3.24) and (3.25), and the arguments $(x,x_{d},u_{d},t)$ for each matrix are suppressed for notational simplicity.
Furthermore, under these equivalent contraction conditions, Theorems LABEL:THM:Thm:Robust_contraction_original and LABEL:THM:Thm:robuststochastic hold for the virtual systems (3.17) and (3.18), respectively. In particular, if $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ of (2.32) holds, or equivalently
@@
where $V_{s\ell}=\int^{x}_{x_{d}}\delta q^{\top}M\delta q$ and $V_{\ell}=\int^{x}_{x_{d}}\|\Theta\delta q\|$ are as given in Theorem LABEL:THM:Thm:path_integral with $M=\Theta^{\top}\Theta$, the disturbance bounds $\bar{d}_{c}$ and $\bar{g}_{c}$ are given in (3.1) and (3.2), respectively, and $C_{C}=\bar{g}_{c}^{2}({2}{\alpha_{G}}^{-1}+1)$. Note that for stochastic systems, the probability that $\|x-x_{d}\|$ is greater than or equal to $\varepsilon\in\mathbb{R}_{>0}$ is given as
@@
Substituting (3.19) into (3.22) gives (3.23). Since $\nu>0$ and $W\succ 0$, multiplying (3.23) by $\nu$ and then by $W$ from both sides preserves matrix definiteness. Also, the resultant inequalities are equivalent to the original ones [lmi, p. 114]. These operations performed on (3.23) yield (3.24). If $\beta=\alpha_{s}>0$ for stochastic systems, applying Schur’s complement lemma [lmi, p. 7] to (3.24) results in the Linear Matrix Inequality (LMI) condition (3.25) in terms of $\bar{W}$ and $\nu$. Therefore, (3.22)–(3.25) are indeed equivalent.
Also, since we have $\|\partial_{\mu}d(\mu,x,t)\|\leq\bar{d}_{c}$ for $d$ in (3.17) and $\|\partial_{\mu}G(\mu,x,t)\|_{F}^{2}\leq\bar{g}_{c}^{2}$ for $G$ in (3.18), the virtual systems in (3.17) and (3.18) clearly satisfy the conditions of Theorems LABEL:THM:Thm:Robust_contraction_original and LABEL:THM:Thm:robuststochastic if it is equipped with (3.22), which is equivalent to (3.23)–(3.25). This implies the exponential bounds (3.27)–(3.29) rewritten using $\chi=\overline{m}/\underline{m}$, following the proofs of Theorems LABEL:THM:Thm:Robust_contraction_original and LABEL:THM:Thm:robuststochastic. $\qed$
@@
Although the conditions (3.22)–(3.25) depend on $(x_{d},u_{d})$, we could also use the SDC formulation with respect to a fixed point mypaperTAC, mypaper in Lemma 3.1 to make them independent of the target trajectory as in the following theorem.
@@
Contraction Conditions with Fixed SDCsThm:fixed_sdc Let $(\bar{x},\bar{u})$ be a fixed point selected arbitrarily in $\mathbb{R}^{n}\times\mathbb{R}^{m}$, e.g., $(\bar{x},\bar{u})=(0,0)$, and let $A(x,t)$ be an SDC matrix constructed with $({s},\bar{s},\bar{u})=(x,\bar{x},\bar{u})$ in Lemma 3.1, i.e.,
@@
Suppose that the contraction metric of Theorem LABEL:THM:Thm:CVSTEM:LMI is designed by $M(x,t)$ with $A$ of (3.30), independently of the target trajectory $(x_{d},u_{d})$, and that the systems (3.1) and (3.2) are controlled by
@@
with such $M(x,t)$, where $R(x,t)\succ 0$ is a weight matrix on $u$. If the function $\phi(x,x_{d},u_{d},t)=A(\varrho,x,t)(x_{d}-\bar{x})+B(x,t)(u_{d}-\bar{u})$ is Lipschitz in $x$ with its Lipschitz constant $\bar{L}$, then Theorem LABEL:THM:Thm:CVSTEM:LMI still holds with $\alpha$ of the conditions (3.22)–(3.25) replaced by $\alpha+\bar{L}\sqrt{\overline{m}/\underline{m}}$. The same argument holds for state estimation of Theorem LABEL:THM:Thm:CVSTEM-LMI-est to be discussed in Sec. 4.2.
@@
where $K(x,t)=R(x,t)^{-1}B(x,t)^{\top}M(x,t)$. Following the proof of Theorem LABEL:THM:Thm:CVSTEM:LMI, the computation of $\dot{V}$, where $V=\delta q^{\top}M(x,t)\delta q$, yields an extra term
@@
As demonstrated in mypaperTAC, we could directly use the extra term $2\delta q^{\top}M({\partial\phi}/{\partial q})\delta q$ of (3.33) in (3.22)–(3.25) without upper-bounding it, although now the conditions of Theorem LABEL:THM:Thm:fixed_sdc depend on $(x,q,t)$ instead of $(x,t)$. Also, the following two inequalities given in mypaperTAC with $\bar{\gamma}=\nu\gamma$, $\gamma\in\mathbb{R}_{\geq 0}$:
@@
are combined as one LMI (3.25) in Theorems LABEL:THM:Thm:CVSTEM:LMI and LABEL:THM:Thm:fixed_sdc.
@@
Contraction in $\mathcal{H}_{\infty}$ Controlex:hinf_lmi The inequalities in Theorem LABEL:THM:Thm:CVSTEM:LMI can be interpreted as in the Riccati inequality in $\mathcal{H}_{\infty}$ control. Consider the following system:
@@
where $A\in\mathbb{R}^{n\times n}$, $B_{u}\in\mathbb{R}^{n\times m}$, $B_{w}\in\mathbb{R}^{n\times w}$, and $C_{z}\in\mathbb{R}^{o\times n}$ are constant matrices, $w\in\mathbb{R}^{w}$ is an exogenous input, and $z\in\mathbb{R}^{o}$ is a system output. As shown in 256331 and [lmi, p. 109], there exists a state feedback gain $K=R^{-1}B_{u}^{\top}P$ such that the $\mathcal{L}_{2}$ gain of the closed-loop system (3.35), $\sup_{\|w\|\neq 0}{\|z\|}/{\|w\|}$, is less than or equal to $\gamma$ if
@@
has a solution $P\succ 0$, where $R\succ 0$ is a constant weight matrix on the input $u$. If we select $B_{w}$ and $C_{z}$ to have $B_{w}B_{w}^{\top}\succeq(P^{-1})^{2}$ and $C_{z}^{\top}C_{z}\succeq 2\alpha P$ for some $\alpha>0$, the contraction condition (3.23) in Theorem LABEL:THM:Thm:CVSTEM:LMI can be satisfied with $M=P$, $B=B_{u}$, and $\beta=1/\gamma^{2}$ due to (3.36).
In Sections 3.3 and 3.4, we will discuss the relationship to input-output stability theory as in Example LABEL:EX:ex:hinf_lmi, using the results of Theorem LABEL:THM:Thm:CVSTEM:LMI.
3.3. Bounded Real Lemma in Contraction Theory
The LMI (3.25) of Theorem LABEL:THM:Thm:CVSTEM:LMI can be interpreted as the LMI condition for the bounded real lemma [brlbook, p. 369]. Let us consider the following Hurwitz linear system with $\varsigma(t)\in\mathcal{L}_{2e}$, (i.e., $\|(\varsigma)_{\tau}\|_{\mathcal{L}_{2}}<\infty$ for $\tau\in\mathbb{R}_{\geq 0}$, see Sec. 1.3):